# A high-order recombination algorithm for weak approximation of stochastic differential equations

### v1

@@ Abstract.
The authors present an algorithm for the application of the high-order recombination method first introduced by Terry Lyons and Christian Litterer in “high-order recombination and an application to cubature on Wiener space” (Ann. Appl. Probab 22(4):1301-1327, 2012), to real-world problems in mathematical finance. They also give numerical examples showing that the algorithm presented successfully avoids the explosive increase in the number of support of the measure that achieves high-order approximations.
@@ 1. Introduction
In [16, 17] they show that the probability measure occurring in weak approximations of a stochastic differential equation (SDE) can be made to avoid an exponential increase in the number of points of its support with respect to accuracy by using a method of transforming the probability measure of a given finite probability space into a probability measure with a smaller support set, while preserving its moments up to a certain degree. This method is the subject of this paper and referred to as high-order recombination method. The aim of this paper is to present a theoretically guaranteed and practically feasible algorithm for the high-order recombination method, and to show that the method is applicable to real mathematical finance problems while maintaining the theoretically predicted performance.
@@ 1.1. Weak approximation problem of SDEs
Let $(\Omega,\mathcal{F},P)$ be a probability space, $\begin{pmatrix}B^{1}(t)&\dots&B^{d}(t)\end{pmatrix}$ be the $d$-dimensional standard Brownian motion, $B^{0}(t)=t$, and $C^{\infty}_{b}(\mathbb{R}^{N};\mathbb{R}^{N})$ be the set of $\mathbb{R}^{N}$-valued smooth functions defined in $\mathbb{R}^{N}$ whose derivatives of any order are bounded. $I_{N}$ denotes the identity map on $\mathbb{R}^{N}$. We also let $X(t,x)$ denote an $\mathbb{R}^{N}$-valued diffusion process defined by the Stratonovich-form SDE:
@@
where $x\in\mathbb{R}^{N}$ and $V_{0},\dots,V_{d}$ are tangent vector fields on $\mathbb{R}^{N}$ whose coefficients belong to $C^{\infty}_{b}(\mathbb{R}^{N};\mathbb{R}^{N})$, that is $V_{i}I_{N}=\begin{pmatrix}(V_{i}I_{N})^{1}&\dots&(V_{i}I_{N})^{N}\end{pmatrix}\in C_{b}^{\infty}(\mathbb{R}^{N};\mathbb{R}^{N})$, and the equality $V_{i}g(y)=\displaystyle{\sum_{j=1}^{N}(V_{i}I_{N})^{j}(x)\frac{\partial g}{\partial x_{j}}(y)}$ holds for all $i\in\{0,1,\dots,d\}$, $g\in C^{\infty}(\mathbb{R}^{N})$ and $y\in\mathbb{R}^{N}$. Furthermore, SDE (1) can be expressed in an alternative form by utilising the Ito integral $\mathrm{d}B^{i}(t)$:
@@
for all $k\in\{1,\dots,N\}$. This equality is called Ito–Stratonovich transformation.
In this paper, we address the weak approximation problem of SDEs, that is the numerical calculation of $E[f\left(X\left(T,x\right)\right)]$, where $X\left(t,x\right)$ is the value of an $N$-dimensional diffusion process at time $t$ and $f$ is a $\mathbb{R}$-valued function defined on $\mathbb{R}^{N}$ with some regularity. This calculation is crucial in many areas of applied science and engineering, for example, pricing and hedging financial derivatives. There are two main approaches to solve this problem numerically: using numerical methods for partial differential equations and performing simulations. Here, we focus on simulations, particularly on so-called tree-based simulation method described in the next subsection.
@@ 1.2. High-order cubature on Wiener space method
We adopt the high-order cubature on Wiener space methods [18, 22] in this paper. In the cubature on Wiener space methods, one approximates the infinite-dimensional Wiener measure with a finite-dimensional measure on a given partition of time $[0,T]$. We refer to this partition as $\Delta=\{0=t_{0}<\dots<t_{n}=T\}$ and, define $\sharp\Delta=n$ and $s_{j}=t_{j}-t_{j-1}$.
@@
As a simple illustrative example of cubature measure on Wiener space, a cubature of degree 3, called cub3, which is proposed in [18, 24] is outlined below.
$\begin{split}X^{({\rm cub3},\Delta)}_{t_{0}}(\eta)&=x,\\ X^{(\rm{cub3},\Delta)}_{\lambda t_{j+1}+(1-\lambda)t_{j}}(\eta)&=\exp\left(\lambda\left(s_{j+1}V_{0}+\sum_{i=1}^{d}\sqrt{s_{j+1}}\eta_{i,j+1}V_{i}\right)\right)X^{(\rm{cub3},\Delta)}_{t_{j}}(\eta)\end{split}$ (4)
@@
for a tangent vector field $V$ on $\mathbb{R}^{N}$ and $\eta=\left\{\eta_{i,j}\right\}_{\begin{subarray}{c}i=1,\dots,d\\ j=1,\dots,\sharp\Delta\end{subarray}}$ denotes a $d\sharp\Delta$-dimensional vector consisting of i.i.d. random variables such that $P(\eta_{i,j}=\pm 1)=1/2$. We remark that for $s\in\mathbb{R}$, $\exp(sV)(x)$ is well-defined by virtue of the chain rule and solves the ODE (5).
Then we obtain a $3$rd-order cubature measure on Wiener space $\mu^{(\rm{cub3},\Delta)}$ supported by $\left\{\omega_{\eta}\left|\,\eta\in\{-1,1\}^{d\sharp\Delta}\right.\right\}$ where
@@
and each $\omega_{\eta}$ has an equal weight of $2^{-d\sharp\Delta}$. Therefore, the computational cost of calculating
@@ 1.2.2. Tree-based simulation
Discretisation methods similar to the cub3 method, which proceeds by drawing a direct product of discrete random variables $\eta$, are called tree based simulations. This terminology stems from the structural interpretation that the support $\left\{\left.\omega_{\eta}\,\right|\,\eta\in\{-1,1\}^{d\sharp\Delta}\right\}$ of the measure can be considered as a tree, with the root $\left.\omega_{\eta}\right|_{s=0}$ and the leaves $\left\{\left.\left.\omega_{\eta}\right|_{s=T}\,\right|\,\eta\in\{-1,1\}^{d\sharp\Delta}\right\}$.
We will prepare some notations for tree-based simulations. Let $\{X^{({\rm Alg},\Delta)}\left(t_{i}\right)\}_{i=0,1,\dots,\sharp\Delta}$ denote a discretised process by a discretisation method Alg using discrete random variables with respect to the partition $\Delta$ of interval $[0,T]$. We denote a probability measure on $\mathbb{R}^{N}$ induced by $X^{({\rm Alg},\Delta)}\left(t_{i}\right)$ as $\mu_{t_{i}}^{({\rm Alg},\Delta)}=\sum_{j=1}^{m_{i}}w_{j}^{(i)}\delta_{p_{j}^{(i)}}$, where $\left\{\left(p_{j}^{(i)},w_{j}^{(i)}\right)\right\}_{j=1\dots,m_{i}}$ is the set of pairs consisting of points $p_{j}^{(i)}$ in $\mathbb{R}^{N}$ and weights $w_{j}^{(i)}$ assigned to them. $\delta_{x}$ denote the Dirac measure at $x\in\mathbb{R}^{N}$. We also denote the transition probability as
@@
Then, we define one step forward operator on the measure space as
@@
Then, we get the measure $\mu_{T}^{(\rm{Alg},\Delta)}=\hat{Q}^{\left({\rm Alg},\Delta\right)}_{(s_{\sharp\Delta})}\circ\cdots\circ\hat{Q}^{\left({\rm Alg},\Delta\right)}_{(s_{1})}\left(\delta_{x}\right)$. The process of approximating $E[f\left(X\left(T,x\right)\right)]$ by the expectation
@@
is referred to as weak approximation by tree-based simulation. The cub 3 method $\mu_{T}^{(\rm{cub3},\Delta)}$, introduced in Subsection 1.2.1, is one of the simplest examples of tree-based simulation methods. The Ninomiya–Victoir method ([24]) $\mu_{T}^{(\rm{NV},\Delta)}$, which serves as the subject of the numerical experiments in this paper, is introduced in Subsection 2.3.
@@ 1.2.3. Order of discretisation
The order of discretisation method is defined as follows.
@@
We call Alg a $p$th order discretisation or discretisation of order $p$ if there exists a positive constant $C_{p}$ and for all $n\in\mathbb{N}$ a partitioning $\Delta$ of $[0,T]$ such that $\sharp\Delta=n$ exists and the following inequality
$\left\lvert E\left[f\left(X(t,x)\right)\right]-E\left[f\left(X^{({\rm Alg},\Delta)}\left(T\right)\right)\right]\right\rvert<C_{p}(\sharp\Delta)^{-p}$ (10)
holds.
In the following, we will refer to discretisation methods of second or higher order as high-order discretisation. Some specific methods of high-order discretisation method based on cubature on Wiener space are proposed in [24, 20, 26], which is introduced later in subsection 2.3. In addition, research on the construction of high-order cubature measures has explored various avenues, including simple algebraic construction [8], randomized construction [9], and Hopf algebras and unshuffle expansions-based construction [6]. For ODEs, a series of lower-order discretisation methods that are combined by allowing negative weights to obtain high-order discretisations are called extrapolations. With regard to the weak approximation of SDEs, the extrapolation of the Euler–Maruyama method is studied in [1]. Furthermore, as demonstrated in [7, 25], extrapolation methods can be integrated with [24] to derive discretisations of any order.
@@ 1.3. High-order recombination
When implementing the high-order cubature measure on Wiener space, one often suffers from the problem of exponential growth in the number of $\operatorname{supp}\left(\mu_{T}^{(\rm{Alg},\Delta)}\right)$ as the number of time partitions $\sharp\Delta$ increases. For example, it can be seen from subsection 1.2.1 that the cardinality of the support of $\mu_{T}^{(\rm{cub3},\Delta)}$ is $2^{d\sharp\Delta}$, while the cardinality of the support of $\mu_{T}^{(\rm{NV},\Delta)}$ is $9^{d\sharp\Delta}$, where NV is the second order discretisation method to be explained in subsection 2.3. In the following, we refer to this problem as the support explosion problem.
At present, the most widely practised methods for overcoming this problem are those of a partial sampling method, such as the Monte Carlo and quasi-Monte Carlo methods. Among the partial sampling methods, the method called TBBA is particularly applicable and effective for the tree-based simulation methods [2, 21, 19]. Furthermore, in [3, 4, 22], the authors successfully apply tree-based simulation with TBBA to the numerical calculations of forward backward SDEs.
High-order recombination is proposed in [17] as an alternative solution to the support explosion problem. There, the authors show that high-order recombination can be used to keep the growth rate of the cardinality of the support of the measure in polynomial order with respect to the required approximation accuracy. Several applications of high-order recombination are discussed in [15]. In [27], a number of algorithms for obtaining reduced measures are discussed. The notion of reduced measure plays a significant role in the construction of recombination measures, and will be introduced in the subsequent subsection 1.3.1.
An overview of the procedure for constructing what we call high-order recombination measures, which are the main subject of this paper, is given in this subsection.
@@ 1.3.1. Reduced measure
We first introduce the concept of a reduced measure, which is a measure with reduced support preserving relations defined by a set of test functions. Let $\mu$ be a discrete probability measure on a space $\tilde{\Omega}$, and $G=\{g_{1},g_{2},\dots,g_{\mathcal{M}_{G}}\}$ be a set of test functions, where $\mathcal{M}_{G}$ is the number of test functions and each $g_{i}:\tilde{\Omega}\rightarrow\mathbb{R}$ is a measurable function.
@@
    $\int g(x)\tilde{\mu}(dx)=\int g(x)\mu(dx)\quad\text{for all}\quad g\in G$,
@@
In [17], they show the existence of a reduced measure for a finite Borel measure $\mu$ on a polish space $\tilde{\Omega}$. In this paper, for the discretisation method of order $m$ and $\tilde{\Omega}=\mathbb{R}^{N}$, we consider the set of test functions $G$ as the distinct homogeneous polynomials of degree at most $m$ and assume that $g_{1}(x)=1$, $g_{2}(0)=g_{3}(0)=\dots=g_{\mathcal{M}_{G}}(0)=0$. Note that the condition (ii) determined by $g_{1}(x)=1$ ensures that the reduced measure $\tilde{\mu}$ also becomes a probability measure.

### v2

@@ Abstract.
This paper presents an algorithm for applying the high-order recombination method, originally introduced by Lyons and Litterer in “High-order recombination and an application to cubature on Wiener space” (Ann. Appl. Probab. 22(4):1301–1327, 2012), to practical problems in mathematical finance. A refined error analysis is provided, yielding a sharper condition for space partitioning. Based on this condition, a computationally feasible recursive partitioning algorithm is developed. Numerical examples are also included, demonstrating that the proposed algorithm effectively avoids the explosive growth in the cardinality of the support required to achieve high-order approximations.
@@ 1. Introduction
In [16, 17], it is shown that the probability measure arising in weak approximations of a stochastic differential equation (SDE) can avoid exponential growth in the cardinality of its support with respect to accuracy, by means of a method that transforms the probability measure of a given finite probability space into one with smaller support, while preserving moments up to a certain degree. This method is the subject of this paper and is referred to as the high-order recombination method. The aim of the paper is to present a theoretically guaranteed and practically feasible algorithm for the high-order recombination method, based on a refined error analysis and a sharp condition for space partitioning, and to demonstrate its applicability to real problems in mathematical finance while maintaining the theoretically predicted performance.
@@ 1.1. Weak approximation problem of SDEs
Let $(\Omega,\mathcal{F},P)$ be a probability space, and let $\begin{pmatrix}B^{1}(t)&\dots&B^{d}(t)\end{pmatrix}$ denote the $d$-dimensional standard Brownian motion. For notational convenience, we define $B^{0}(t)=t$. We denote by $C^{\infty}_{b}(\mathbb{R}^{N};\mathbb{R}^{N})$ the set of $\mathbb{R}^{N}$-valued smooth functions on $\mathbb{R}^{N}$ whose derivatives of all orders are bounded. Let $I_{N}$ denote the identity map on $\mathbb{R}^{N}$. Let $X(t,x)$ be an $\mathbb{R}^{N}$-valued diffusion process defined as the solution to the Stratonovich-form SDE:
@@
where $x\in\mathbb{R}^{N}$ and $V_{0},\dots,V_{d}$ are tangent vector fields on $\mathbb{R}^{N}$, each having coefficients in $C^{\infty}_{b}(\mathbb{R}^{N};\mathbb{R}^{N})$; that is, $V_{i}I_{N}=\begin{pmatrix}(V_{i}I_{N})^{1}&\dots&(V_{i}I_{N})^{N}\end{pmatrix}\in C_{b}^{\infty}(\mathbb{R}^{N};\mathbb{R}^{N})$, and the identity $V_{i}g(y)=\displaystyle{\sum_{j=1}^{N}(V_{i}I_{N})^{j}(x)\frac{\partial g}{\partial x_{j}}(y)}$ holds for all $i\in\{0,1,\dots,d\}$, $g\in C^{\infty}(\mathbb{R}^{N})$ and $y\in\mathbb{R}^{N}$. Furthermore, the SDE (1) can be equivalently written in Itô form:
@@
for all $k\in\{1,\dots,N\}$. This identity is known as the Itô–Stratonovich transformation.
In this paper, we address the weak approximation problem for SDEs; that is, the numerical evaluation of $E[f\left(X\left(T,x\right)\right)]$, where $X\left(t,x\right)$ denotes the value of an $N$-dimensional diffusion process at time $t$ and $f$ is a $\mathbb{R}$-valued function defined on $\mathbb{R}^{N}$ with appropriate regularity. This type of computation is crucial in many areas of applied science and engineering—for example, in the pricing and hedging of financial derivatives. There are two main approaches to solving this problem numerically: one based on partial differential equations and the other on simulations. We focus on the latter approach, particularly on a so-called tree-based simulation method, which is described in a later subsection.
@@ 1.2. High-order cubature on Wiener space method
This section adopts high-order cubature on Wiener space methods [18, 22]. In these methods, one approximates the infinite-dimensional Wiener measure with a finite-dimensional measure on a given partition of the time interval $[0,T]$. We denote this partition by $\Delta=\{0=t_{0}<\dots<t_{n}=T\}$, and define $\sharp\Delta=n$ and $s_{j}=t_{j}-t_{j-1}$.
@@
As a simple illustrative example of a cubature measure on Wiener space, we outline a degree 3 cubature, called cub3, which is proposed in [18, 24].
$\begin{split}X^{({\rm cub3},\Delta)}_{t_{0}}(\eta)&=x,\\ X^{(\rm{cub3},\Delta)}_{\lambda t_{j+1}+(1-\lambda)t_{j}}(\eta)&=\\ \exp&\left(\lambda\left(s_{j+1}V_{0}+\sum_{i=1}^{d}\sqrt{s_{j+1}}\eta_{i,j+1}V_{i}\right)\right)X^{(\rm{cub3},\Delta)}_{t_{j}}(\eta)\end{split}$ (4)
@@
for a tangent vector field $V$ on $\mathbb{R}^{N}$, and $\eta=\left\{\eta_{i,j}\right\}_{\begin{subarray}{c}i=1,\dots,d\\ j=1,\dots,\sharp\Delta\end{subarray}}$ denotes a $d\sharp\Delta$-dimensional vector consisting of independent and identically distributed (i.i.d.) random variables satisfying $P(\eta_{i,j}=\pm 1)=1/2$. We also remark that for any $s\in\mathbb{R}$, the expression $\exp(sV)(x)$ is well-defined by virtue of the chain rule and solves the ODE (5).
We then obtain a $3$rd-order cubature measure on Wiener space $\mu^{(\rm{cub3},\Delta)}$ supported on $\left\{\omega_{\eta}\left|\,\eta\in\{-1,1\}^{d\sharp\Delta}\right.\right\}$ where
@@
and each path $\omega_{\eta}$ has an equal weight of $2^{-d\sharp\Delta}$. Therefore, the computational cost of evaluating
@@ 1.2.2. Tree-based simulation
Discretisation methods similar to the cub3 method, which proceed by drawing a direct product of discrete random variables $\eta$, are referred to as tree based simulations. This terminology stems from the structural interpretation that the support $\left\{\left.\omega_{\eta}\,\right|\,\eta\in\{-1,1\}^{d\sharp\Delta}\right\}$ of the measure can be viewed as a tree, with the root $\left.\omega_{\eta}\right|_{s=0}$ and the leaves $\left\{\left.\left.\omega_{\eta}\right|_{s=T}\,\right|\,\eta\in\{-1,1\}^{d\sharp\Delta}\right\}$.
We now introduce notation for tree-based simulations. Let $\{X^{({\rm Alg},\Delta)}\left(t_{i}\right)\}_{i=0}^{\sharp\Delta}$ denote the discretised process obtained by a discretisation method Alg, based on discrete random variables with respect to the partition $\Delta$ of the interval $[0,T]$. We denote by $\mu_{t_{i}}^{({\rm Alg},\Delta)}$ the probability measure on $\mathbb{R}^{N}$ induced by $X^{({\rm Alg},\Delta)}(t_{i})$ and write it as
$\mu_{t_{i}}^{({\rm Alg},\Delta)}=\sum_{j=1}^{m_{i}}w_{j}^{(i)}\delta_{p_{j}^{(i)}},$ (6)
where $\left\{\left(p_{j}^{(i)},w_{j}^{(i)}\right)\right\}_{j=1}^{m_{i}}$ is the set of pairs consisting of points $p_{j}^{(i)}\in\mathbb{R}^{N}$ and the corresponding weights $w_{j}^{(i)}$. Here, $\delta_{x}$ denotes the Dirac measure at $x\in\mathbb{R}^{N}$. We also denote the transition probability as
@@
We then define the one-step forward operator on the measure space as
@@
Finally, we obtain the measure $\mu_{T}^{(\rm{Alg},\Delta)}=\hat{Q}^{\left({\rm Alg},\Delta\right)}_{(s_{\sharp\Delta})}\circ\cdots\circ\hat{Q}^{\left({\rm Alg},\Delta\right)}_{(s_{1})}\left(\delta_{x}\right)$. The process of approximating $E[f\left(X\left(T,x\right)\right)]$ by the expectation
@@
is referred to as a weak approximation by tree-based simulation. The cub3 method $\mu_{T}^{(\rm{cub3},\Delta)}$, introduced in subsection 1.2.1, is one of the simplest examples of tree-based simulation methods. The Ninomiya–Victoir method ([24]) $\mu_{T}^{(\rm{NV},\Delta)}$, which serves as the subject of the numerical experiments in this paper, is introduced in subsection 2.3.
@@ 1.2.3. Order of discretisation
The order of a discretisation method is defined as follows.
@@
We call Alg a discretisation of order $p$ or a $p$th order discretisation if there exists a constant $C_{p}>0$ such that for all $n\in\mathbb{N}$ there exists a partition $\Delta$ of $[0,T]$ with $\sharp\Delta=n$ satisfying the inequality
$\left\lvert E\left[f\left(X(t,x)\right)\right]-E\left[f\left(X^{({\rm Alg},\Delta)}\left(T\right)\right)\right]\right\rvert<C_{p}(\sharp\Delta)^{-p}.$ (11)
In what follows, we refer to discretisation methods of order $2$ or higher as high-order discretisation methods. Some specific high-order discretisation method based on cubature on Wiener space are proposed in [24, 20, 26], and the one used in this paper is introduced in subsection 2.3. In addition, research on the construction of high-order cubature measures has explored various avenues, including simple algebraic construction [8], randomised construction [9], and constructions based on Hopf algebras and unshuffle expansions [6]. For ODEs, discretisation methods of fixed order when linearly combined using weights that may include negative values can yield methods of higher order. This approach is commonly referred to as extrapolation. In the context of weak approximation for SDEs, the extrapolation of the Euler–Maruyama method is studied in [1]. Furthermore, as demonstrated in [7, 25], extrapolation methods can be combined with the method of [24] to construct discretisations of arbitrary order.
@@ 1.3. High-order recombination
When implementing high-order cubature measures on Wiener space, one often encounters the problem of exponential growth in the cardinality of the $\operatorname{supp}\left(\mu_{T}^{(\rm{Alg},\Delta)}\right)$ as the number of time partitions $\sharp\Delta$ increases. For example, as seen in subsection 1.2.1 the cardinality of the support of $\mu_{T}^{(\rm{cub3},\Delta)}$ is $2^{d\sharp\Delta}$ and that of the support of $\mu_{T}^{(\rm{NV},\Delta)}$ is $9^{d\sharp\Delta}$, where NV refers to the second-order discretisation method introduced in subsection 2.3. In what follows, we refer to this issue as the support explosion problem.
At present the most widely used approach to overcome this problem is partial sampling such as Monte Carlo and quasi-Monte Carlo methods. Among these the technique known as TBBA is particularly effective and well suited to the tree-based simulation methods [2, 21, 19]. Furthermore in [3, 4, 22], tree-based simulation with TBBA is successfully applied to the numerical calculation of forward–backward SDEs.
High-order recombination is proposed in [17] as an alternative approach to resolving the support explosion problem. There the authors show that the cardinality of the support can be kept polynomial in the required approximation accuracy. Several applications of high-order recombination are presented in [15]. In [27] a number of algorithms for obtaining reduced measures are discussed. The notion of reduced measure plays a significant role in the construction of recombination measures, and will be introduced in the subsequent subsection 1.3.1.
This subsection presents an overview of the procedure for constructing what we call high-order recombination measures, which form the main subject of this paper.
@@ 1.3.1. Reduced measure
We first introduce the concept of a reduced measure, which is a measure with reduced support preserving relations defined by a set of test functions. Let $\mu$ be a discrete probability measure on a space $\tilde{\Omega}$, and let $G=\{g_{1},g_{2},\dots,g_{\mathcal{M}_{G}}\}$ be a set of test functions, where $\mathcal{M}_{G}$ denotes the number of test functions and each $g_{i}:\tilde{\Omega}\rightarrow\mathbb{R}$ is a measurable function.
@@
    $\int g(x)\,\tilde{\mu}(\mathrm{d}x)=\int g(x)\,\mu(\mathrm{d}x)\quad\text{for all}\quad g\in G$,
@@
In [17], the existence of a reduced measure is shown for a finite Borel measure $\mu$ on a Polish space $\tilde{\Omega}$.
In this paper for discretisation methods of order $m$ and with $\tilde{\Omega}=\mathbb{R}^{N}$, we consider the set of test functions $G$ to consist of distinct homogeneous polynomials of degree at most $m$ and we assume that $g_{1}(x)=1$ and $g_{2}(0)=g_{3}(0)=\dots=g_{\mathcal{M}_{G}}(0)=0$.
Note that condition (ii), determined by $g_{1}(x)=1$, ensures that the reduced measure $\tilde{\mu}$ is also a probability measure.